OpenStudy (anonymous):

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. 2x2 + 6x – 7 = 0

OpenStudy (anonymous):

\[2x ^{2} + 6x - 7?\]

OpenStudy (kinggeorge):

Do you know what the discriminant is?

OpenStudy (anonymous):

Yeah, sorry just clarifying what the function is.

OpenStudy (kinggeorge):

Do both these names belong to the same person? Anyways, calculate the discriminant. In this the discriminant is \(6^2-4(2)(-7)=36+56=92\) Since this is positive, it must have two real roots. However, since 92 isn't a perfect square, the two roots must be irrational.